OpenStudy (anonymous):

Find the area of a trapezoid if the altitude is 6" and the median is 8." _ sq. in.

OpenStudy (anonymous):

48 sq in

OpenStudy (anonymous):

area of trapezoid is median \( \times \) height

OpenStudy (anonymous):

Ahh, thank you, I over complicated it in my head when I was looking at the equation the whole time :P hahaha, thanks.

OpenStudy (anonymous):

Glad to help.

OpenStudy (anonymous):

because\[\frac{(a+c)}{2}=median\] area is\[\frac{(a+c)}{2}*h\]

OpenStudy (anonymous):